OpenStudy (anonymous):

Indicate in standard form the equation of the line passing through the given points. S(1/2, 1), T(1/2, 4)

OpenStudy (anonymous):

yay. y=mx+b BUTTTTT find the slope first. Sooo: x2-x1/y2-y1. (1/2-1/2)/1-4 slope=0. using y=mx+b, Substitute one of the pair of coordinates in, and see what you get! y=mx+b 1=(0)(1/2)+b 1=b. y=1. der ya go. =D

OpenStudy (anonymous):

what does x=?

OpenStudy (anonymous):

it could equal the x coordinate of one of the pair of coordinates given.